# CCL-LGS: Contrastive Codebook Learning for 3D Language Gaussian Splatting

### v1

@@ 1 Introduction
This task is particularly challenging due to factors such as semantic ambiguity (e.g., a point on the tip of the nose may correspond to both “nose” and “face” queries) and the long-tail distribution of vocabulary queries. Early methods like LERF [8] tackle these issues by extending NeRF [16] with multi-scale CLIP features and dynamic 2D mapping selection, achieving accurate semantic localization. However, their reliance on exhaustive multi-scale rendering leads to inefficiency, and patch-based feature extraction often fails to capture precise object boundaries, resulting scale misalignment and performance degradation. Building upon the rise of explicit 3D representations such as 3DGS [7], recent works [20, 22, 18] differ from earlier NeRF-based methods by leveraging visual foundation models to extract dense, pixel-level semantic features that guide the construction of 3D semantic fields. While differing in implementation, these methods generally follow a common paradigm: supervising 3D semantic representations by projecting them to 2D views and comparing the rendered results with features extracted from pre-trained vision-language models.
@@ 3.2 Two-Level Semantic Feature Extraction
In our method, a uniform 32$\times$32 point prompt is provided to SAM to generate three types of masks corresponding to the semantic scales of subparts, parts, and whole objects. Recognizing that different points may yield conflicting scale assignments (for example, a subpart for one point might be considered a part for another), we merge the subpart and part masks as well as the whole and part masks to obtain two aggregated sets of masks, denoted as $M^{sp}_{o}$ and $M^{wp}_{o}$. A filtering procedure is then applied to eliminate redundant masks based on predicted IoU scores, stability metrics, and overlap rates, with the final non-overlapping masks $M_{i}^{l}$ (where $l\in\{sp,wp\}$ and $i$ indexes the masks) selected based on the product of the IoU and stability scores. For clarity, we omit the superscript $l$ in all subsequent expressions (e.g., $M_{i}$ instead of $M_{i}^{l}$), with the implicit understanding that operations are independently applied to each scale $l$. After obtaining segmentation maps, pixel-aligned language embeddings can be extracted using CLIP. For each pixel $v$, its semantic feature $F_{i}(v)$ can be expressed as:
@@ 3.3 Contrastive Codebook Learning
The first step corresponds to the “Mask Association” module at the top of Fig. 2. Specifically, we first propagate the $K$ masks from the first frame to all frames using SAM2 [24]. For the $t$-th frame, the propagated $K$ masks are compared with all masks $M_{i}$ in that frame using IoU. If the maximum IoU between a propagated mask and a mask $M_{i}$ exceeds 0.5, then $M_{i}$ is assigned the category of the matching propagated mask; otherwise, it is assigned the category $-1$ to indicate an unmatched mask. Hence, each mask $M_{i}$ is assigned a label $y_{i}\in\{1,2,\dots,K,-1\}$.
@@
Specifically, for each pixel $v$, its semantic feature $F_{i}(v)$ is obtained via Eq. (3). The index $j^{*}$ is assigned to $v$ through Eq. (4), generating a semantic index map $\mathcal{M}\in\mathbb{R}^{H\times W}$, where $H$ and $W$ denote the image height and width, respectively. Subsequently, a feature map $\hat{F}\in\mathbb{R}^{H\times W\times d_{f}}$ is rendered via rasterization and alpha blending (see Sec. 3.1). This feature map is processed by a lightweight MLP decoder followed by a softmax layer to produce a semantic feature distribution ^$\hat{\mathcal{M}}\in\mathbb{R}^{H\times W\times N}$, where $N$ corresponds to the number of codebook categories. To jointly optimize the semantic features of 3D Gaussians and the parameters of the MLP decoder, we minimize the cross-entropy loss:
$\mathcal{L}_{\text{CE}}=\text{CE}(\hat{\mathcal{M}},\mathcal{M}),$ (9)
@@ 4.1 Experiments on LERF
To validate the effectiveness of the proposed Contrastive Codebook Learning (CCL) module, we conduct a series of experiments, including visual analysis of 2D supervision features and ablation studies on 3D semantic segmentation. As shown in Fig. 3, features refined by the CCL module exhibit enhanced spatial precision and stronger activation on semantically relevant regions, eliminating the scattered and ambiguous responses observed in the baseline. This visual evidence demonstrates that CCL improves the alignment between extracted features and semantic ground truth during 2D supervision generation.
We further perform ablation studies to quantitatively analyze the impact of individual loss components on 3D semantic understanding. Evaluations are conducted on four scenes from the LERF dataset. The baseline model uses codebook-based feature compression to produce 2D semantic supervision for 3D segmentation. We compare three variants: baseline with pull loss, baseline with push loss, and the full model with both losses. As shown in Tab. 2, both losses are essential for optimal performance—removing either leads to noticeable degradation, though all variants still outperform the baseline.
Qualitative results in Fig. 5 further reveal the complementary roles of the two losses. The pull loss enhances intra-class consistency, particularly for occluded or partially visible objects like the “glass of water”, while the push loss suppresses false activations in confusing regions, such as around the “kamaboko”. The full model combines both effects to ensure robust and discriminative 3D semantic segmentation across challenging scenes.
@@ 4.2 Experiments on 3D-OVS
We present our qualitative experimental results in Figure 6. Since LEGaussians is a patch-based method, it learns over-smoothed features and fails to capture sharp object boundaries. LangSplat adopts a three-scale architecture but does not enhance multi-view features as our method does. In our experiments, we observed that it sometimes selects inappropriate scales for certain objects, leading to suboptimal segmentation performance. Among the compared methods, our method produces the most accurate segmentation maps, further demonstrating the effectiveness of our CCL-LGS.
@@ 5 Conclusion
In this paper, we introduced CCL-LGS, a novel method for constructing 3D language fields that enables accurate and efficient open-vocabulary queries in 3D spaces. Our work tackles a previously overlooked challenge: applying CLIP directly to imperfect masks produces inconsistent semantic features, and using these features to supervise 3D generation can lead to artifacts. To address this issue, we proposed a dedicated CCL module that constructs compact and distinct features, thereby providing more reliable supervision. Extensive experiments demonstrate that our approach achieves state-of-the-art performance. Our method is limited by the capabilities of SAM and CLIP, with performance affected by imperfect instance masks, and future work will focus on improving mask refinement for greater robustness.

### v2

@@ 1 Introduction
This task is particularly challenging due to factors such as semantic ambiguity (e.g., a point on the tip of the nose may correspond to both “nose” and “face” queries) and the long-tail distribution of vocabulary queries. Early methods like LERF [8] tackle these issues by extending NeRF [16] with multi-scale CLIP features and dynamic 2D mapping selection, achieving accurate semantic localization. However, their reliance on exhaustive multi-scale rendering leads to inefficiency, and patch-based feature extraction often fails to capture precise object boundaries, resulting in scale misalignment and performance degradation. Building upon the rise of explicit 3D representations such as 3DGS [7], recent works [20, 22, 18] differ from earlier NeRF-based methods by leveraging visual foundation models to extract dense, pixel-level semantic features that guide the construction of 3D semantic fields. While differing in implementation, these methods generally follow a common paradigm: supervising 3D semantic representations by projecting them to 2D views and comparing the rendered results with features extracted from pre-trained vision-language models.
@@ 3.2 Two-Level Semantic Feature Extraction
In our method, a uniform $\times32$ point prompt is provided to SAM to generate three types of masks corresponding to the semantic scales of subparts, parts, and whole objects. Recognizing that different points may yield conflicting scale assignments (for example, a subpart for one point might be considered a part for another), we merge the subpart and part masks as well as the whole and part masks to obtain two aggregated sets of masks, denoted as $M^{sp}_{o}$ and $M^{wp}_{o}$. A filtering procedure is then applied to eliminate redundant masks based on predicted IoU scores, stability metrics, and overlap rates, with the final non-overlapping masks $M_{i}^{l}$ (where $l\in\{sp,wp\}$ and $i$ indexes the masks) selected based on the product of the IoU and stability scores. For clarity, we omit the superscript $l$ in all subsequent expressions (e.g., $M_{i}$ instead of $M_{i}^{l}$), with the implicit understanding that operations are independently applied to each scale $l$. After obtaining segmentation maps, pixel-aligned language embeddings can be extracted using CLIP. For each pixel $v$, its semantic feature $F_{i}(v)$ can be expressed as:
@@ 3.3 Contrastive Codebook Learning
The first step corresponds to the “Mask Association” module at the top of Fig. 2. Specifically, we first propagate the $K$ masks from the first frame to all frames using SAM2 [24]. For the $t$-th frame, the propagated $K$ masks are compared with all masks $M_{i}$ in that frame using IoU. If the maximum IoU between a propagated mask and a mask $M_{i}$ exceeds 0.5, then $M_{i}$ is assigned the category of the matching propagated mask; otherwise, it is assigned the category 1-1 to indicate an unmatched mask. Hence, each mask $M_{i}$ is assigned a label $y_{i}\in\{1,2,\dots,K,-1\}$.
@@
Specifically, for each pixel $v$, its semantic feature $F_{i}(v)$ is obtained via Eq. (3). The index $j^{*}$ is assigned to $v$ through Eq. (4), generating a semantic index map $\mathcal{M}\in\mathbb{R}^{H\times W}$, where $H$ and $W$ denote the image height and width, respectively. Subsequently, a feature map $\hat{F}\in\mathbb{R}^{H\times W\times d_{f}}$ is rendered via rasterization and alpha blending (see Sec. 3.1). This feature map is processed by a lightweight MLP decoder followed by a softmax layer to produce a semantic feature distribution $\hat{\mathcal{M}}\in\mathbb{R}^{H\times W\times N}$, where $N$ corresponds to the number of codebook categories. To jointly optimize the semantic features of 3D Gaussians and the parameters of the MLP decoder, we minimize the cross-entropy loss:
[…]
@@ 4.1 Experiments on LERF
To validate the effectiveness of our Contrastive Codebook Learning (CCL) module, we conduct experiments, including visual analysis of 2D supervision features and ablation studies on 3D semantic segmentation. As shown in Fig. 3, features refined by CCL exhibit higher spatial precision and stronger activation on semantically relevant regions, eliminating scattered and ambiguous responses in the baseline. This demonstrates that CCL improves alignment between extracted features and semantic ground truth during 2D supervision.
We further perform ablation studies to analyze the impact of individual loss components on 3D semantic understanding. Evaluations are conducted on four LERF scenes. The baseline uses codebook-based feature compression to produce 2D semantic supervision for 3D segmentation. We compare three variants: baseline with pull loss, baseline with push loss, and the full model with both. As shown in Tab. 2, both losses are essential for optimal performance—removing either causes noticeable degradation, though all variants still surpass the baseline.
Qualitative results in Fig. 5 further show the complementary roles of the two losses. The pull loss improves intra-class consistency, especially for occluded or partly visible objects like the “glass of water”, while the push loss reduces false activations in confusing regions, such as near the “kamaboko”. Combining both ensures robust, discriminative 3D semantic segmentation in challenging scenes.
@@ 4.2 Experiments on 3D-OVS
We present our qualitative experimental results in Figure 6. Since LEGaussians is a patch-based method, it learns over-smoothed features and fails to capture sharp object boundaries. LangSplat adopts a three-scale architecture but does not enhance multiview features as our method does. In our experiments, we observed that it sometimes selects inappropriate scales for certain objects, leading to suboptimal segmentation performance. Among the compared methods, our method produces the most accurate segmentation maps, further demonstrating the effectiveness of our CCL-LGS.
@@ 5 Conclusion
In this paper, we present CCL-LGS, a novel method to build 3D language fields enabling accurate, efficient open-vocabulary queries. We address an overlooked challenge: directly applying CLIP to imperfect masks yields inconsistent semantic features, introducing artifacts in 3D reconstruction. To solve this, we propose a dedicated CCL module that creates compact, distinct features for more reliable supervision. Extensive experiments show state-of-the-art performance. Limitations remain due to inherent capabilities of SAM and SAM2, as imperfect masks still affect results. Future work will refine masks for greater robustness.